OpenStudy (anonymous):

In circle A, the radius is 7 and the length ofarc C Tis 21. Determine thethe measure of angle C A T.

OpenStudy (anonymous):

if you could please help i can give medals

OpenStudy (anonymous):

so the P = 21+2*7

OpenStudy (anonymous):

http://clip2net.com/s/5DibMd

OpenStudy (anonymous):

ok so P=35?

OpenStudy (anonymous):

after we find the perimeter we use the other formula and substract P=r( angle+2)

OpenStudy (anonymous):

35=7(angle +2)

OpenStudy (anonymous):

\[\theta=5-2=3\]

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Circular_sector

OpenStudy (anonymous):

ok thank you very much! how do i give you a medal?

OpenStudy (anonymous):

click on best response

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

yw